06 February 2008

RHIC pp Set-Up:  After abandoning the new working point, reloaded the old magnet settings from Run06.  Circulating beam in Blue and Yellow by 5:30, with a first ramp by 7:10.  Store dispersion looked good by 4:30 AM.  A 12 bunch ramp was performed with cogging in an attempt to see collisions.  There were no good collision signals yet.  We will try for collisions again tonight. 

An AGS Cooling failure took the beam from 6:30 - 9:00.

Polarimeter measures emittances of 50p in both rings.



Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020